### v2

@@ 4 Experiments and Results
Figure 3: Accuracy Curves for NiNformer Architecture
@@
Figure 4: Loss Curves for NiNformer Architecture
@@ 5 Conclusion
In this work, a newly designed Network in Network block was introduced that substitutes the Attention block used traditionally by Transformer architectures. By extending the token mixing approach presented in the MLP-Mixer to function as a gating signal generator and taking advantage of the gating mechanism to introduce dynamic behavior in the newly proposed block, thus enhancing the static weight approach of the MLP-Mixer by utilizing its layers as a sub-unit network within the outer network formulation, The experimental results show that in comparison to baseline architectures that were chosen from different families of Transformer formulations, our proposed block significantly outperforms, showing noticeable improvements on those baselines, specifically showing a great enhancement of accuracy compared to the standalone MLP-Mixer architecture that acts as a sub-unit, which validates that our proposal, with its dynamic gating of the upstream representation, properly enhances and circumvents the shortcoming of the static weight approach of the standalone MLP-Mixer while still providing more simplicity of operations in contrast to the vanilla ViT Transformer architecture. Future directions of this work are to investigate a multitude of sub-unit network selections, aiming for further enhancements and capabilities.

### v3

@@ 4 Results
Figure 3: An illustration of the accuracy curves for NiNformer architecture.
@@
Figure 4: An illustration of the loss curves for NiNformer architecture.
@@ 5 Conclusion
We introduced in this work a newly designed Network in Network block that substitutes the attention block used traditionally in transformer architectures by extending the token mixing approach presented in the MLP-Mixer to function as a gating signal generator and taking advantage of the gating mechanism to introduce dynamic behavior in the newly proposed block, thus enhancing the static weight approach of the MLP-Mixer by utilizing its layers as a sub-unit network within the outer network formulation. The experimental results show that in comparison to baseline architectures that were chosen from different families of transformer formulations, our proposed block significantly outperforms, showing noticeable improvements on those baselines, specifically showing a great enhancement of accuracy compared to the standalone MLP-Mixer architecture that acts as a sub-unit, which validates that our proposal, with its dynamic gating of the upstream representation, properly enhances and circumvents the shortcoming of the static weight approach of the standalone MLP-Mixer while still providing more simplicity of operations in contrast to the vanilla ViT transformer architecture. Future directions of this work are to investigate a multitude of sub-unit network selections, aiming for further enhancements and capabilities.